# On the Use of Agentic Coding: An Empirical Study of Pull Requests on GitHub

### v2

@@ 4.3. $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent?
Approach. We examined all 475 merged Agentic-PRs and identified 214 Agentic-PRs (45.145.1%) that had modifications between initial submission and merge. We performed manual classification on these 214 Agentic-PRs to understand what types of changes were made during the review process. We manually inspect all commits added after the initial submission and classify them using the same two-dimensional framework from Zeng et al. (DBLP:conf/icse/ZengZQL25) described in Section 4.1. For each PR, we aggregated revision types at the PR level to avoid over-representing PRs with numerous small commits. For example, if a PR included three bug fixes and two stylistic changes across separate commits, it was counted as one instance of each change type. This manual classification was performed by the second and third authors. As in Section 4.1, we treated this as a multi-label task and measured agreement at the label level. The overall agreement rate was 64.9% (541 out of 844 labels). Disagreements were resolved by the second author following the same process described in Section 4.1. In addition, to understand the extent of continued AI assistance during the revision process, we analyzed co-authorship patterns in revision commits. We extracted commit metadata to identify commits containing “Co-Authored-By: Claude” attribution.
@@
fix Code changes that fix bugs and faults within the codebase 45.145.1%
@@
docs Updates to documentation or comments, such as README edits, typo fixes, or API docs improvements 27.427.4%
@@
refactor Code restructuring without changing its behavior, aiming to improve maintainability 25.725.7%
@@
style Non-functional code changes that improve readability or consistency. This type encompasses aspects like variable naming, indentation, and addressing linting or code analysis warnings 22.122.1%
@@
chore Project‐wide housekeeping tasks such as dependency bumps, version increments, and miscellaneous cleanup 19.919.9%
@@
test Additions or modifications to test files, including new test cases or updates to existing tests 15.515.5%
@@
feat Code changes that introduce new features to the codebase, encompassing both internal and user-oriented features 14.614.6%
@@
build Changes to build configurations (e.g., Maven, Gradle, Cargo). Change examples include updating dependencies, configuring build configurations, and adding scripts 13.313.3%
ci Changes to CI/CD workflows and configurations, e.g., “.travis.yml” and “.github/workflows” 6.66.6%
@@
perf Code changes that improve performance, such as enhancing execution speed or reducing memory consumption 1.31.3%
@@
Results. The majority (45.145.1%) of revisions to Agentic-PRs target bug fixes. As shown in Table 4, the most common revision type was fixing functional bugs, accounting for 45.145.1% (102 out of 226) of revised Agentic-PRs. These fixes typically involved critical error handling improvements, and representative revisions addressed situations where critical errors were improperly handled. In one case, [^36] https://github.com/qltysh/qlty/pull/1591/commits/87444f54e702a64f519ca1052380bf9f50f7be96 concurrent error propagation required introducing channel-based communication mechanisms to ensure fatal errors were immediately surfaced from worker closures. Similarly, another revision [^37] https://github.com/qltysh/qlty/pull/1588/commits/a38e75cc15657d8864cff66b388189b1ea677e19 upgraded file operation failures from warnings to fatal errors, recognizing that certain failure conditions should halt execution rather than proceed with a potentially corrupted state. These patterns reveal that agent-generated code frequently implements optimistic error handling strategies that fail to distinguish between recoverable and non-recoverable failure conditions.
@@
Addressing documentation gaps constitutes the second largest focus (27.427.4%) of Agentic-PRs revisions. Documentation updates were present in 27.427.4% (62 out of 226) of revisions, as shown in Table 4. Although agents sometimes generated or updated code comments, they often failed to synchronize all relevant artifacts. For example, one revision [^38] https://github.com/brentyi/tyro/pull/266/commits/8d3bf92a48a8bef4dbb09bfc31bc39dbfb544046 removed an outdated optional dependencies section from installation documentation that should have been deleted alongside related code changes. This disconnect means that, in practice, human reviewers spend a non-trivial amount of effort ensuring that documentation, README files, and code comments accurately reflect code changes from agents.
@@
Refactoring is frequently required (25.725.7%) to improve code structure before merging Agentic-PRs. Table 4 shows that 58 out of 226 modified Agentic-PRs involved refactoring, highlighting the need for structural improvements in these submissions. These structural changes, which do not affect program behaviour, were frequently necessary to eliminate code duplication, enhance modularity, or revise architectural decisions made by the agent. For example, a reviewer consolidated redundant initialization logic spread across multiple entry points into a unified service. [^39] https://github.com/basicmachines-co/basic-memory/pull/88/commits/ca5fa557c25bead9aafcf545544957d835bba798 In this case, common functionality for database migrations and file synchronization was extracted from separate components and moved into a shared module, improving both error handling and logging consistency. This finding suggests that although the initial submissions from Agentic-PRs were functionally correct, reviewers often had to refactor them to align better with project architecture, enhance maintainability, and reduce technical debt.
@@
A common portion (22.122.1%) of revisions is commonly focused on polishing code style rather than behaviour. As shown in Table 4, 50 out of 226 modified Agentic-PRs received style-related revisions during review, indicating code quality issues in initial submissions. Typical fixes include enforcing naming conventions, correcting formatting, and resolving linter warnings that agentic tools missed. For example, a reviewer addressed multiple static analysis violations that were present in the agent’s original code, [^40] https://github.com/INCATools/ontology-access-kit/pull/831/commits/5a6f71986fcfaafa291b9aa405d51d76c8c27c97 including unused imports, exception variables that were declared but never referenced, and incorrect import ordering. These changes are largely cosmetic but necessary for integration, and suggest that current agents often underperform on project-specific style rules, requiring human intervention to maintain readability and consistency. Static analysis violations and ignored best practices were common triggers for these types of revisions.
@@
Project Housekeeping tasks account for 19.919.9% of revisions. According to Table 4, chores represent 19.919.9% (45 out of 226) of revisions, typically involving project-wide metadata such as version bumps or release notes that the agent overlooked. For example, one revision [^41] https://github.com/yumemi-inc/Tart/pull/106/commits/88594d8a116e7c91390740cf7e9ddb7ccfaa61f5 involved a simple but critical version bump from “3.0.0-alpha01” to “3.0.0-alpha02” in a Gradle configuration file (libs.versions.toml), a necessary step for a new release that the agent did not perform. This pattern indicates a specific gap in agent capabilities: while they successfully modify application code, they often fail to propagate corresponding changes to project-level configuration files. Consequently, these essential administrative tasks fall to human reviewers to ensure the project’s versioning and release process remains consistent and accurate.
@@
New features are added in 14.614.6% of Agentic-PRs by human reviewers. Table 4 indicates that 33 out of 226 modified Agentic-PRs (14.614.6%) received feature additions during the review process. For example, one revision [^42] https://github.com/github/github-mcp-server/pull/118/commits/0527bc552591df807ee7be0b3a79f428d2f0ed2c expanded the PR review API functionality that the agent had initially implemented with basic capabilities. The revision added support for multi-line comments, including new parameters for line positioning, validation logic for parameter combinations, and comprehensive test coverage. This enhancement demonstrates that agent-generated implementations often provide core functionality but miss advanced features or edge cases that developers identify during review, requiring subsequent additions to meet complete user requirements.
@@
Build configuration adjustments occur in 13.313.3% of reviewed Agentic-PRs. Table 4 shows that 30 out of 226 modified Agentic-PRs received build configuration updates during review. One revision [^43] https://github.com/coder/coder/pull/17035/commits/3c3aa219b895a981dd13a1b530603aa257f5549e addressed compatibility issues between the linting tool and the newer Go version. The agent’s initial implementation did not account for version incompatibilities, causing build failures with Go 1.24.1. The revision updated the linter version, configured it to run with a reduced set of checks, and modified the build process to continue despite linting failures. This temporary solution allowed the build pipeline to function while a comprehensive fix was planned for a subsequent update. The revision illustrates that agent-generated build configurations may overlook toolchain compatibility requirements, requiring manual intervention to maintain functional build processes across version updates.
@@
Human reviewers also shore up test coverage that agents frequently omit. Although tests account for only 15.515.5% (35 out of 226) of the revisions in Table 4, the added suites are often substantial and guard edge cases and failure paths that the original submission did not address. For instance, one revision [^44] https://github.com/aleph-im/aleph-vm/pull/786/commits/bf69485b01c6d246126bd471f5f1509eef4cc84d added comprehensive unit tests for GPU X-VGA support detection functionality that were absent from the agent’s initial submission. The new tests cover device-argument generation, detection-process validation, error-handling scenarios, and configuration-flow verification. The agent had implemented the GPU support feature without corresponding test coverage, which reviewers identified as a gap requiring remediation. This pattern echoes our earlier finding in Section 4.2 that reviewers explicitly reject Agentic-PRs when they cannot fully trust the code without further human verification.
@@
CI/CD and performance optimizations represent a smaller but crucial tail of revisions (7.9%). CI/CD modifications are relatively rare, comprising 6.66.6% of revisions, yet they are essential for maintaining a healthy pipeline. One revision, for instance, introduced a script to enforce version-string consistency across build artifacts after the original agent submission broke continuous integration. [^45] https://github.com/coder/coder/pull/17164/commits/de29681960767cb7eb01b407fa6de4fa48a07976 Performance improvements appear in an even smaller fraction, just 1.31.3%, such as when a reviewer added a cache-busting mechanism to ensure data freshness for an efficient but potentially stale storage driver implemented by an agent, thereby achieving both accuracy and performance. [^46] https://github.com/giselles-ai/giselle/pull/695/commits/f9356079fe49ce56ca04e821098432916367032b Taken together, these results highlight that while agentic coding already addresses the majority of functional and stylistic changes, human reviewers continue to play a key role in ensuring project hygiene, pipeline reliability, and targeted performance optimization.
@@
Answer to RQ4 Among merged Agentic PRs, 45.145.1% require reviewer revisions, primarily for bug fixes (45.145.1%), documentation updates (27.427.4%), refactoring (25.725.7%), and code style improvements (22.122.1%). This highlights that while AI-generated code is a strong starting point, human oversight is essential to ensure correctness, maintainability, and adherence to project standards.
@@ 5.1. Implications for researchers
Create PR-centric benchmarks that reward alignment across code, documentation, and review resolution. Current benchmarks such as SWE-bench (DBLP:conf/iclr/JimenezYWYPPN24) overlook documentation drift, style adherence, and reviewer negotiation, all of which accounted for 48.5% of Agentic-PR revisions (Section 4.4). This gap between benchmark evaluation and real-world requirements represents a fundamental misalignment in how we assess coding agent capabilities. While existing benchmarks focus predominantly on functional correctness, whether the generated code passes test cases, they fail to capture nearly half of the actual work required to integrate AI-generated code into production systems. We suggest building datasets that incorporate time-to-merge, rework size, and the alignment between documentation and implementation. Such benchmarks can be applied during both training and evaluation of coding agents to promote improvements beyond functional correctness.
Instrument the entire human-AI workflow, not isolated tools. Our results show that 41.6% of revision commits were co-authored by Claude Code (Section 4.4), and review threads frequently referenced GitHub Copilot and other agents. This widespread multi-agent collaboration highlights a critical gap in our understanding of agent-assisted development workflows. We recommend that future studies collect comprehensive full-trace datasets that capture the entire development lifecycle from initial prompts and agent planning phases to intermediate code generations, CI/CD logs, review iterations, and final merge decisions. Such datasets would reveal crucial interaction patterns, including where hand-offs between agents and developers occur, where duplicate efforts waste resources, and where conflicting edits from different agents inflate latency and introduce defects.

### v3

@@ 4.3. $RQ_{3}$: What proportion of Agentic-PRs are accepted without revisions? If needed, to what extent?
Approach. We examined all 475 merged Agentic-PRs and identified 214 Agentic-PRs (45.145.1%) that had modifications between initial submission and merge. We performed manual classification on these 214 Agentic-PRs to understand what types of changes were made during the review process. We manually inspect all commits added after the initial submission and classify them using the same two-dimensional framework from Zeng et al. (Zeng et al., 2025) described in Section 4.1. For each PR, we aggregated revision types at the PR level to avoid over-representing PRs with numerous small commits. For example, if a PR included three bug fixes and two stylistic changes across separate commits, it was counted as one instance of each change type. This manual classification was performed by the second and third authors. As in Section 4.1, we treated this as a multi-label task and measured agreement at the label level. The overall agreement rate was 64.1% (541 out of 844 labels). Disagreements were resolved by the second author following the same process described in Section 4.1. In addition, to understand the extent of continued AI assistance during the revision process, we analyzed co-authorship patterns in revision commits. We extracted commit metadata to identify commits containing “Co-Authored-By: Claude” attribution.
@@
fix Code changes that fix bugs and faults within the codebase 47.747.7%
@@
docs Updates to documentation or comments, such as README edits, typo fixes, or API docs improvements 29.029.0%
@@
refactor Code restructuring without changing its behavior, aiming to improve maintainability 27.127.1%
@@
style Non-functional code changes that improve readability or consistency. This type encompasses aspects like variable naming, indentation, and addressing linting or code analysis warnings 23.423.4%
@@
chore Project‐wide housekeeping tasks such as dependency bumps, version increments, and miscellaneous cleanup 21.021.0%
@@
test Additions or modifications to test files, including new test cases or updates to existing tests 16.416.4%
@@
feat Code changes that introduce new features to the codebase, encompassing both internal and user-oriented features 15.415.4%
@@
build Changes to build configurations (e.g., Maven, Gradle, Cargo). Change examples include updating dependencies, configuring build configurations, and adding scripts 14.014.0%
ci Changes to CI/CD workflows and configurations, e.g., “.travis.yml” and “.github/workflows” 7.07.0%
@@
perf Code changes that improve performance, such as enhancing execution speed or reducing memory consumption 1.41.4%
@@
Results. The majority (47.747.7%) of revisions to Agentic-PRs target bug fixes. As shown in Table 4, the most common revision type was fixing functional bugs, accounting for 47.747.7% (102 out of 214) of revised Agentic-PRs. These fixes typically involved critical error handling improvements, and representative revisions addressed situations where critical errors were improperly handled. In one case, [^36] https://github.com/qltysh/qlty/pull/1591/commits/87444f54e702a64f519ca1052380bf9f50f7be96 concurrent error propagation required introducing channel-based communication mechanisms to ensure fatal errors were immediately surfaced from worker closures. Similarly, another revision [^37] https://github.com/qltysh/qlty/pull/1588/commits/a38e75cc15657d8864cff66b388189b1ea677e19 upgraded file operation failures from warnings to fatal errors, recognizing that certain failure conditions should halt execution rather than proceed with a potentially corrupted state. These patterns reveal that agent-generated code frequently implements optimistic error handling strategies that fail to distinguish between recoverable and non-recoverable failure conditions.
@@
Addressing documentation gaps constitutes the second largest focus (29.029.0%) of Agentic-PRs revisions. Documentation updates were present in 29.029.0% (62 out of 214) of revisions, as shown in Table 4. Although agents sometimes generated or updated code comments, they often failed to synchronize all relevant artifacts. For example, one revision [^38] https://github.com/brentyi/tyro/pull/266/commits/8d3bf92a48a8bef4dbb09bfc31bc39dbfb544046 removed an outdated optional dependencies section from installation documentation that should have been deleted alongside related code changes. This disconnect means that, in practice, human reviewers spend a non-trivial amount of effort ensuring that documentation, README files, and code comments accurately reflect code changes from agents.
@@
Refactoring is frequently required (27.127.1%) to improve code structure before merging Agentic-PRs. Table 4 shows that 58 out of 214 modified Agentic-PRs involved refactoring, highlighting the need for structural improvements in these submissions. These structural changes, which do not affect program behaviour, were frequently necessary to eliminate code duplication, enhance modularity, or revise architectural decisions made by the agent. For example, a reviewer consolidated redundant initialization logic spread across multiple entry points into a unified service. [^39] https://github.com/basicmachines-co/basic-memory/pull/88/commits/ca5fa557c25bead9aafcf545544957d835bba798 In this case, common functionality for database migrations and file synchronization was extracted from separate components and moved into a shared module, improving both error handling and logging consistency. This finding suggests that although the initial submissions from Agentic-PRs were functionally correct, reviewers often had to refactor them to align better with project architecture, enhance maintainability, and reduce technical debt.
@@
A common portion (23.423.4%) of revisions is commonly focused on polishing code style rather than behaviour. As shown in Table 4, 50 out of 214 modified Agentic-PRs received style-related revisions during review, indicating code quality issues in initial submissions. Typical fixes include enforcing naming conventions, correcting formatting, and resolving linter warnings that agentic tools missed. For example, a reviewer addressed multiple static analysis violations that were present in the agent’s original code, [^40] https://github.com/INCATools/ontology-access-kit/pull/831/commits/5a6f71986fcfaafa291b9aa405d51d76c8c27c97 including unused imports, exception variables that were declared but never referenced, and incorrect import ordering. These changes are largely cosmetic but necessary for integration, and suggest that current agents often underperform on project-specific style rules, requiring human intervention to maintain readability and consistency. Static analysis violations and ignored best practices were common triggers for these types of revisions.
@@
Project Housekeeping tasks account for 21.021.0% of revisions. According to Table 4, chores represent 21.021.0% (45 out of 214) of revisions, typically involving project-wide metadata such as version bumps or release notes that the agent overlooked. For example, one revision [^41] https://github.com/yumemi-inc/Tart/pull/106/commits/88594d8a116e7c91390740cf7e9ddb7ccfaa61f5 involved a simple but critical version bump from “3.0.0-alpha01” to “3.0.0-alpha02” in a Gradle configuration file (libs.versions.toml), a necessary step for a new release that the agent did not perform. This pattern indicates a specific gap in agent capabilities: while they successfully modify application code, they often fail to propagate corresponding changes to project-level configuration files. Consequently, these essential administrative tasks fall to human reviewers to ensure the project’s versioning and release process remains consistent and accurate.
@@
New features are added in 15.415.4% of Agentic-PRs by human reviewers. Table 4 indicates that 33 out of 214 modified Agentic-PRs (15.415.4%) received feature additions during the review process. For example, one revision [^42] https://github.com/github/github-mcp-server/pull/118/commits/0527bc552591df807ee7be0b3a79f428d2f0ed2c expanded the PR review API functionality that the agent had initially implemented with basic capabilities. The revision added support for multi-line comments, including new parameters for line positioning, validation logic for parameter combinations, and comprehensive test coverage. This enhancement demonstrates that agent-generated implementations often provide core functionality but miss advanced features or edge cases that developers identify during review, requiring subsequent additions to meet complete user requirements.
@@
Build configuration adjustments occur in 14.014.0% of reviewed Agentic-PRs. Table 4 shows that 30 out of 214 modified Agentic-PRs received build configuration updates during review. One revision [^43] https://github.com/coder/coder/pull/17035/commits/3c3aa219b895a981dd13a1b530603aa257f5549e addressed compatibility issues between the linting tool and the newer Go version. The agent’s initial implementation did not account for version incompatibilities, causing build failures with Go 1.24.1. The revision updated the linter version, configured it to run with a reduced set of checks, and modified the build process to continue despite linting failures. This temporary solution allowed the build pipeline to function while a comprehensive fix was planned for a subsequent update. The revision illustrates that agent-generated build configurations may overlook toolchain compatibility requirements, requiring manual intervention to maintain functional build processes across version updates.
@@
Human reviewers also shore up test coverage that agents frequently omit. Although tests account for only 16.416.4% (35 out of 214) of the revisions in Table 4, the added suites are often substantial and guard edge cases and failure paths that the original submission did not address. For instance, one revision [^44] https://github.com/aleph-im/aleph-vm/pull/786/commits/bf69485b01c6d246126bd471f5f1509eef4cc84d added comprehensive unit tests for GPU X-VGA support detection functionality that were absent from the agent’s initial submission. The new tests cover device-argument generation, detection-process validation, error-handling scenarios, and configuration-flow verification. The agent had implemented the GPU support feature without corresponding test coverage, which reviewers identified as a gap requiring remediation. This pattern echoes our earlier finding in Section 4.2 that reviewers explicitly reject Agentic-PRs when they cannot fully trust the code without further human verification.
@@
CI/CD and performance optimizations represent a smaller but crucial tail of revisions (8.4%). CI/CD modifications are relatively rare, comprising 7.07.0% of revisions, yet they are essential for maintaining a healthy pipeline. One revision, for instance, introduced a script to enforce version-string consistency across build artifacts after the original agent submission broke continuous integration. [^45] https://github.com/coder/coder/pull/17164/commits/de29681960767cb7eb01b407fa6de4fa48a07976 Performance improvements appear in an even smaller fraction, just 1.41.4%, such as when a reviewer added a cache-busting mechanism to ensure data freshness for an efficient but potentially stale storage driver implemented by an agent, thereby achieving both accuracy and performance. [^46] https://github.com/giselles-ai/giselle/pull/695/commits/f9356079fe49ce56ca04e821098432916367032b Taken together, these results highlight that while agentic coding already addresses the majority of functional and stylistic changes, human reviewers continue to play a key role in ensuring project hygiene, pipeline reliability, and targeted performance optimization.
@@
Answer to RQ4 Among merged Agentic PRs, 45.145.1% require reviewer revisions, primarily for bug fixes (47.747.7%), documentation updates (29.029.0%), refactoring (27.127.1%), and code style improvements (23.423.4%). This highlights that while AI-generated code is a strong starting point, human oversight is essential to ensure correctness, maintainability, and adherence to project standards.
@@ 5.1. Implications for researchers
Create PR-centric benchmarks that reward alignment across code, documentation, and review resolution. Current benchmarks such as SWE-bench (Jimenez et al., 2024) overlook documentation drift, style adherence, and reviewer negotiation, all of which accounted for 52.4% of Agentic-PR revisions (Section 4.4). This gap between benchmark evaluation and real-world requirements represents a fundamental misalignment in how we assess coding agent capabilities. While existing benchmarks focus predominantly on functional correctness, whether the generated code passes test cases, they fail to capture nearly half of the actual work required to integrate AI-generated code into production systems. We suggest building datasets that incorporate time-to-merge, rework size, and the alignment between documentation and implementation. Such benchmarks can be applied during both training and evaluation of coding agents to promote improvements beyond functional correctness.
Instrument the entire human-AI workflow, not isolated tools. Our results show that 34.1% of revision commits were co-authored by Claude Code (Section 4.4), and review threads frequently referenced GitHub Copilot and other agents. This widespread multi-agent collaboration highlights a critical gap in our understanding of agent-assisted development workflows. We recommend that future studies collect comprehensive full-trace datasets that capture the entire development lifecycle from initial prompts and agent planning phases to intermediate code generations, CI/CD logs, review iterations, and final merge decisions. Such datasets would reveal crucial interaction patterns, including where hand-offs between agents and developers occur, where duplicate efforts waste resources, and where conflicting edits from different agents inflate latency and introduce defects.